OpenStudy (anonymous):

if f(x)=2x-1 and g(x) =cosx find (g composite f)(x)

OpenStudy (anonymous):

you want: \[g(f(x)) = g(2x-1) = \cos(2x-1)\]

OpenStudy (anonymous):

yes